# Post-selection inference for quantifying uncertainty in changes in variance

### v1

@@ 1 Introduction
In this paper we consider a related approach, of post-selection inference [24]. The idea is to construct a $p$-value for each detected changepoint, with the these $p$-values being valid in finite samples, albeit under strong distributional assumptions for the data. This is a challenging problem in general as, if we were to use the data to estimate changepoint locations, naively re-using the same data to estimate the uncertainty associated with each changepoint will lead to biased results. The idea of post-selection inference is, when testing for the presence of a changepoint at some detected location $\tau$, we condition on the information in the data that led to detection at $\tau$ when calculating the $p$-value. To date, almost all work in this area has focused on the univariate change in mean case, e.g. [18, 20, 9, 6], or the piecewise constant polynomial model [27]. As far as we know, there are no current approaches for quantifying the evidence for detected changes in variance. In this paper we fill this gap by extending the post-selection inference ideas to cover the change in variance setting.
@@
The post-selection inference methods for the change in mean problem use the fact that the natural test statistic for a change at a given location will have a standard normal distribution under the null. To account for the fact that we have used the same data to detect changes and thus to decide which locations to test, we need to condition this distribution on the fact that the tested location is a detected changepoint. This leads to a conditional null distribution that is constrained to lie on an interval [18] or a union of intervals [20] of the real line – depending on precisely what information is conditioned on. Importantly it is straightforward to calculate this interval or intervals for methods for detecting a change in mean that use the popular CUSUM statistic within a binary segmentation algorithm or within a penalised likelihood approach. One of the advantages of this approach is that it is general: it applies to a wide range of methods for detecting changes, and also to most natural null hypotheses that one may wish to test.
@@
In this paper, we consider the two most common ways of detecting a change in variance. The first uses the idea that a change in variance in some time series, $X_{t}$ say, will correspond to a change in mean in $X_{t}^{2}$. Thus we can shift the data so it has mean zero, and apply a method for detecting a change in mean to the square of this shifted data. The second method is where we estimate changepoint locations by directly maximizing the likelihood of the piecewise constant variance model. In both cases we then use the same test statistic for testing whether a detected change is real. This test statistic is equivalent to the likelihood-ratio test statistic for a change in variance in Gaussian data, and is the proportion of the sum of squares of the data in the region prior to the putative change. If we had used separate information, or independent data, to choose which putative change location to test then the null distribution of this test statistic a Beta distribution with known parameters. For our setting we need to calculate the null distribution conditional on detecting the change, which leads to constraining the Beta distribution to a union of intervals of $[0,1]$.
@@
To calculate the post-selection $p$-values requires us to calculate these intervals. For the first approach to detecting changes, this can be done analytically using the same ideas as for the change in mean setting. For the second approach, this is no longer possible so we show how Monte Carlo can be used to approximate the $p$-values [see 36, for similar use of Monte Carlo in post-selection inference].
@@ 2 Testing for a change in variance
$\phi=\frac{C_{(\hat{\tau}-h+1):\hat{\tau}}^{2}}{C_{(\hat{\tau}-h+1):(\hat{\tau}+h)}^{2}}.$
Under $H_{0}$, $\phi\sim\text{Beta}\left(\frac{h}{2},\frac{h}{2}\right)$, with $E(\phi)=\frac{1}{2}$. Under $H_{1}$, we would expect $\phi$ to be closer to 00 if the variance increases after $\hat{\tau}$, and closer to $1$ if $\hat{\tau}$ is followed by a decrease in variance.
@@ 2.1 Post-selection inference
The $p$-value in (3) is valid if we use separate information, or independent data, to calculate the $p$-value from that which we use to choose the location to test. However, if we use the same data to calculate the $p$-value as we have used to estimate changepoint locations, this introduces bias into the $p$-value. In order to get valid $p$-values in this case, we need to calculate the probability of (3) and conditional on the information used to choose to test for a change at $\hat{\tau}$: here, this is the fact that $\hat{\tau}$ is in the set of estimated changepoints. Letting $\mathcal{M}(\boldsymbol{X})$ denote the set of estimated changepoints $\{\hat{\tau}_{1},\ldots,\hat{\tau}_{k}\}$ which we obtain when we apply a particular changepoint algorithm to $\boldsymbol{X}$, we therefore want to calculate the $p$-value in (3) conditional on the fact that $\hat{\tau}\in\mathcal{M}(\boldsymbol{X})$:
@@
Our null hypothesis (2) is that the variance is constant inside the region $\{\hat{\tau}-h+1,\ldots,\hat{\tau}+h\}$. This does not specify the value that $\sigma_{t}^{2}$ takes for any $t$ outside this window, nor the value of the fixed variance within the window. Therefore, in order for the test to be well-defined, we must condition on sufficient statistics for these parameters: the values of $X_{t}$ outside of $\{\hat{\tau}-h+1,\ldots,\hat{\tau}+h\}$, and the sum of squares within the window: $C_{(\hat{\tau}-h+1):(\hat{\tau}+h)}^{2}=\sum_{t=\hat{\tau}-h+1}^{\hat{\tau}+h}X_{t}^{2}$.
For now, we will also condition on the parameters within the window which are orthogonal to $\phi$, following the approach of [20] for the change in mean case. This makes the $p$-values more straightforward to compute, whilst still resulting in a valid test (i.e. uniform $p$-values under $H_{0}$), although it may result in some loss of power [17, 20]. Later on, in Section 5.3, we consider relaxing this condition.
@@
$\begin{split}W_{i}^{l}&=\frac{C_{(\hat{\tau}-h+1):(\hat{\tau}-h+i)}^{2}}{C_{(\hat{\tau}-h+1):(\hat{\tau}-h+i+1)}^{2}},~{}~{}i=1,\ldots,h-1;\\ W_{i}^{r}&=\frac{C_{(\hat{\tau}+1):(\hat{\tau}+1)}^{2}}{C_{(\hat{\tau}+1):(\hat{\tau}+i+1)}^{2}},~{}~{}i=1,\ldots,h-1.\end{split}$
The following results shows how we can write the square of $\{X_{\hat{\tau}-h+1},\ldots,X_{\hat{\tau}+h}\}$ in terms of $\phi$, $C^{2}_{0}$, and $W^{r}_{k}$ and $W^{l}_{k}$ for $k=1,\ldots,h-1$, and that the latter set of variables are independent under the null hypothesis.
@@
Therefore, conditioning on the dimensions of the data which are orthogonal to $\phi$ is equivalent to conditioning on the values of $W_{1}^{l},\ldots,W_{h-1}^{l}$, $W_{1}^{r},\ldots,W_{h-1}^{r}$, and $C_{0}^{2}$. We can see from (4) above that doing this while allowing only $\phi$ to change has the effect of multiplying $X_{\hat{\tau}-h+1,\ldots,X_{\hat{\tau}}}$ by $\sqrt{\frac{\phi}{\phi_{obs}}}$ and multiplying $X_{\hat{\tau}+1},\ldots,X_{\hat{\tau}+h}$ by $\sqrt{\frac{1-\phi}{1-\phi_{obs}}}$. (This transformation is illustrated in Figure 1.) Hence, let
@@
We consider only the subset of possible data sets $\{\boldsymbol{X}^{\prime}(\phi):\phi\in\mathbb{R}\}$. Letting $\mathcal{S}$ denote the set $\mathcal{S}=\{\phi:\hat{\tau}\in\mathcal{M}(\boldsymbol{X}^{\prime}(\phi))\}$, the post-selection $p$-value we want to calculate is
@@ 3.1 CUSUM for change in variance
$G_{s,e}(t)=\sqrt{\frac{(t-s+1)(e-t)}{e-s+1}}\left(\bar{Y}_{s:t}-\bar{Y}_{(t+1):e}\right),~{}~{}t=s,\ldots,e-1.$
@@ 3.1.1 Calculating the post-selection $p$-value
To calculate $\mathcal{S}$ we can borrow ideas from methods used for the change in mean model. [18] showed that by conditioning on the set of changes $\mathcal{M}(\boldsymbol{X})$, the order in which the changes are found, and the direction of each change (positive or negative), we can write down a set of inequalities which must be satisfied by the CUSUM statistic. For a given segment $(s,e)$, in order to detect a change point at $\hat{\tau}$ we must have $|G_{s,e}(\hat{\tau})|\geq\lambda$, where $\lambda$ is the changepoint detection threshold, and
@@
$\boldsymbol{X}^{\prime}(\phi)=\boldsymbol{X}_{obs}-\frac{\boldsymbol{\nu}}{||\boldsymbol{\nu}||_{2}^{2}}\phi_{obs}+\frac{\boldsymbol{\nu}}{||\boldsymbol{\nu}||_{2}^{2}}\phi.$
Since $G_{s,e}(t)$ is a linear function of $\boldsymbol{X}$, and $\boldsymbol{X}^{\prime}(\phi)$ is a linear function of $\phi$, the CUSUM statistic in terms of $\phi$ is also linear in $\phi$. Hence, we can write the set of inequalities in $G_{s,e}(t)$ as linear inequalities in $\phi$, and they will therefore hold for an interval of $\phi$ values, which it is straightforward to calculate. [20] shows that we can increase power by extending this idea to avoid conditioning on the order and directions of the changes, by calculating the intervals associated with each possible combination of changes which yield $\hat{\tau}\in\mathcal{M}(\boldsymbol{X})$. This gives us the set $\mathcal{S}$ which is the union of all such intervals.
@@
$\begin{split}G_{s,e}(t)&=\left(\frac{(t-s+1)(e-t)}{e-s+1}\right)^{1/2}\left(\bar{Y}_{s:t}-\bar{Y}_{(t+1):e}\right)\\ &=\left(\frac{(t-s+1)(e-t)}{e-s+1}\right)^{1/2}\left(\frac{1}{t-s+1}\sum_{j=s}^{t}X_{j}^{2}-\frac{1}{e-t}\sum_{j=t+1}^{e}X_{j}^{2}\right),\end{split}$
@@
Using the definition of $\boldsymbol{X}^{\prime}(\phi)$ from (5), we can write each element of this as $X_{j}^{\prime}(\phi)=a_{j}+b_{j}\phi$ where
$a_{j}=\begin{cases}X_{j}^{2}&\phi\leq\hat{\tau}-h\text{ or }\phi>\hat{\tau}+h\\ 0&\hat{\tau}-h<\phi\leq\hat{\tau}\\ \frac{1}{1-\phi_{obs}}X_{j}^{2}&\hat{\tau}<\phi\leq\hat{\tau}+h\end{cases}$
$b_{j}=\begin{cases}0&\phi\leq\hat{\tau}-h\text{ or }\phi>\hat{\tau}+h\\ \frac{1}{\phi_{obs}}X_{j}^{2}&\hat{\tau}-h<\phi\leq\hat{\tau}\\ -\frac{1}{1-\phi_{obs}}X_{j}^{2}&\hat{\tau}<\phi\leq\hat{\tau}+h\end{cases}$
@@ 3.2 Likelihood-ratio based detection of changes
Table 1 demonstrates how this leads to increased accuracy compared with using the CUSUM statistic. We simulated data from a model with $T=400$ and 3 changepoints at $\tau=100,200,300$; the variances on the four segments being $\{1,4,0.25,1\}$, and for each simulated data set we used binary segmentation to estimate three changepoint locations, using both the CUSUM statistic and the likelihood ratio statistic to determine a change. The table records the proportion of runs in which a changepoint was detected within 10 of each true changepoint. We see that CUSUM does reasonably well at detecting the first two changes, but does very badly at locating the third, smaller change. Using the likelihood ratio statistic, however, all three changepoints are detected in over $90\%$ of cases.
@@ 3.2.1 Post-selection inference
where $\mathcal{S}=\{\phi:\hat{\tau}\in\mathcal{M}(\boldsymbol{X}^{\prime}(\phi))\}$. Previously, when detecting changes based on the CUSUM statistic, we were able to write down a set of linear inequalities in $\phi$ which defined the set $\mathcal{S}$. However, while we can write down inequalities in $\Lambda_{s,e}(\phi,\tau)$ as we did for $G_{s,e}(\phi,\tau)$, $\Lambda_{s,e}(\phi,\tau)$ is not a linear function of $\phi$: we can show that
$\Lambda_{s,e}(\tau,\phi)=\log\left(\frac{(A_{s,e}\phi+B_{s,e})^{s-e+1}}{(A_{s,\tau}\phi+B_{s,\tau})^{\tau-s+1}((A_{s,e}-A_{s,\tau})\phi+B_{s,e}-B_{s,\tau})^{e-\tau}}\right),$ (9)
where $A_{s,e}$ and $B_{s,e}$ are constants depending on $s$, $e$, $h$, and $\phi_{obs}$ (see the Appendix for more details). In general, it is not possible to analytically solve these inequalities for $\phi$ as it was earlier.
@@
One way to reduce the variance of $\hat{p}$, without increasing the computation time, is to use stratified sampling. For $i=1,\ldots,N$, we sample $z^{(i)}\sim U(\frac{i-1}{N},\frac{i}{N})$. We then calculate $\phi^{(i)}$ to satisfy $\Pr(\phi<\phi^{(i)})=z^{(i)}$. (Section C.2 in the supplementary material shows how the variance of $\hat{p}$ decreases using stratified, compared to random sampling.) Doing this ensures that our set of sampled $\phi$ values are spread evenly across the sample space, which can help to give more accurate estimates. However, it is not infrequently the case that $\Pr(\phi\in\mathcal{S})$ is of the order of $\frac{1}{N}$ or smaller, in which case using stratified sampling is unlikely to significantly improve the estimate $\hat{p}$, as we will still have very few $\phi^{(i)}\in\mathcal{S}$.
To overcome this we resort to using simulation to estimate the intervals that contribute to $\mathcal{S}$. These estimates can then be used to directly estimate the post-selection $p$-value or to construct better proposal distributions for an importance sampling based Monte Carlo approach.

### v2

@@ 1 Introduction
In this paper we consider post-selection inference methods [3, 26] for quantifying the degree of uncertainty associated with detected changepoints. The idea is to construct a $p$-value for each detected changepoint, with these $p$-values being valid in finite samples, albeit under strong distributional assumptions for the data. This is a challenging problem in general as, if we were to use the data to estimate changepoint locations, naively re-using the same data to estimate the uncertainty associated with each changepoint will lead to biased results. The idea of post-selection inference is, when testing for the presence of a changepoint at some detected location $\tau$, we condition on the information in the data that led to detection at $\tau$ when calculating the $p$-value. To date, almost all work in this area has focused on the univariate change in mean case, e.g. [20, 22, 10, 7], or the piecewise constant polynomial model [29]. As far as we know, there are no current approaches for quantifying the evidence for detected changes in variance. In this paper we fill this gap by extending the post-selection inference ideas to cover the change in variance setting.
@@
The post-selection inference methods for the change in mean problem use the fact that the natural test statistic for a change at a given location will have a normal distribution under the null. To account for the fact that we are using the same data to calculate the test statistic as to detect changes – and thus to decide which locations to test – we need to condition this distribution on the fact that the tested location is a detected changepoint. This leads to a conditional null distribution that is constrained to lie on an interval [20] or a union of intervals [22] of the real line – depending on precisely what information is conditioned on. Importantly it is straightforward to calculate this interval or intervals for methods for detecting a change in mean that use the popular CUSUM statistic within a binary segmentation algorithm or a penalised likelihood approach. One of the advantages of this approach is that it is general: it applies to a wide range of methods for detecting changes, and also to most natural null hypotheses that one may wish to test.
@@
In this paper, we consider the two most common ways of detecting a change in variance. The first uses the idea that a change in variance in some time series, $X_{t}$ say, will correspond to a change in mean in $X_{t}^{2}$. Thus we can shift the data so it has mean zero, and apply a method for detecting a change in mean to the square of this shifted data. The second method is where we estimate changepoint locations by directly maximizing the likelihood of the piecewise constant variance model. In both cases we then use the same test statistic for testing whether a detected change is real. This test statistic is equivalent to the likelihood-ratio test statistic for a change in variance in Gaussian data, and is the proportion of the sum of squares of the data in the region prior to the putative change. If we had used separate information, or independent data, to choose which putative change location to test then the null distribution of this test statistic is a Beta distribution with known parameters. For our setting we need to calculate the null distribution conditional on detecting the change, which leads to constraining the Beta distribution to a union of intervals of $[0,1]$.
@@
To calculate the post-selection $p$-values requires us to calculate these intervals. For the first approach to detecting changes, this can be done analytically using the same ideas as for the change in mean setting. For the second approach, this is no longer possible so we show how Monte Carlo can be used to approximate the $p$-values [see 38, for similar use of Monte Carlo in post-selection inference].
@@ 2 Testing for a change in variance
$\phi=\frac{C_{(\hat{\tau}-h+1):\hat{\tau}}^{2}}{C_{(\hat{\tau}-h+1):(\hat{\tau}+h)}^{2}}=\frac{C_{(\hat{\tau}-h+1):\hat{\tau}}^{2}}{C_{(\hat{\tau}-h+1):\hat{\tau}}^{2}+C_{(\hat{\tau}+1):(\hat{\tau}+h)}^{2}}.$
The terms in the last expression, $C_{(\hat{\tau}-h+1):\hat{\tau}}^{2}$ and $C_{(\hat{\tau}+1):(\hat{\tau}+h)}^{2}$, are independent gamma random variables with shape parameter $h/2$ and scale parameter $2\sigma^{2}$ under $H_{0}$. Thus, under $H_{0}$, $\phi\sim\text{Beta}\left(\frac{h}{2},\frac{h}{2}\right)$. Under $H_{1}$, we would expect $\phi$ to be closer to 00 if the variance increases after $\hat{\tau}$, and closer to $1$ if $\hat{\tau}$ is followed by a decrease in variance.
@@ 2.1 Post-selection inference
The $p$-value in (3) is valid if we use separate information, or independent data, to calculate the $p$-value from that which we use to choose the location to test. However, if we use the same data to calculate the $p$-value as we have used to estimate changepoint locations, this introduces bias into the $p$-value. In particular, as demonstrated by Figure 1, we tend to obtain $p$-values that are much smaller than they should be, so the Type I error rate is too high; this is because in general we have chosen to test only at the locations where the test statistic is close to either 0 or 1, resulting in $p$-values that are too small. In order to get valid $p$-values in this case, we need to calculate the probability of (3) conditional on the information used to choose to test for a change at $\hat{\tau}$: here, this is the fact that $\hat{\tau}$ is in the set of estimated changepoints. Letting $\mathcal{M}(\boldsymbol{X})$ denote the set of estimated changepoints $\{\hat{\tau}_{1},\ldots,\hat{\tau}_{K}\}$ which we obtain when we apply a particular changepoint algorithm to $\boldsymbol{X}$, we therefore want to calculate the $p$-value in (3) conditional on the fact that $\hat{\tau}\in\mathcal{M}(\boldsymbol{X})$:
@@
Figure 1: QQ plot of $p$-values, calculated using (3), for data simulated under the null hypothesis of no change, where we selected change locations to test from running binary segmentation on the same data. Different coloured lines correspond to different choices of $h$ in the null hypothesis. Re-using the data results in $p$-values that are much smaller than they should be.
Our null hypothesis (2) is that the variance is constant inside the region $\{\hat{\tau}-h+1,\ldots,\hat{\tau}+h\}$. This does not specify the value that $\sigma_{t}^{2}$ takes for any $t$ outside this window, nor the value of the fixed variance within the window. Therefore, in order for the test to be well-defined, i.e. have the same distribution for all data generating processes consistent with the null hypothesis and the model, we must condition on sufficient statistics for these parameters: the values of $X_{t}$ outside of $\{\hat{\tau}-h+1,\ldots,\hat{\tau}+h\}$, and the sum of squares within the window: $C_{(\hat{\tau}-h+1):(\hat{\tau}+h)}^{2}=\sum_{t=\hat{\tau}-h+1}^{\hat{\tau}+h}X_{t}^{2}$.
To make calculating the $p$-values easier it is common to additionally condition on the data within the region we are testing that is orthogonal to $\phi$. We can condition on extra information like this whilst still retaining valid $p$-values (i.e. uniform $p$-values under $H_{0}$), but it may result in some loss of power [19, 22]. In Section 5.3, we consider reducing this conditioning.
@@
$\begin{split}W_{i}^{l}&=\frac{C_{(\hat{\tau}-h+1):(\hat{\tau}-h+i)}^{2}}{C_{(\hat{\tau}-h+1):(\hat{\tau}-h+i+1)}^{2}},~{}~{}i=1,\ldots,h-1;\\ W_{i}^{r}&=\frac{C_{(\hat{\tau}+1):(\hat{\tau}+i)}^{2}}{C_{(\hat{\tau}+1):(\hat{\tau}+i+1)}^{2}},~{}~{}i=1,\ldots,h-1.\end{split}$
The following results show how we can write the square of $\{X_{\hat{\tau}-h+1},\ldots,X_{\hat{\tau}+h}\}$ in terms of $\phi$, $C^{2}_{0}$, and $W^{r}_{i}$ and $W^{l}_{i}$ for $i=1,\ldots,h-1$, and that the latter set of variables are independent under the null hypothesis.
@@
Therefore, conditioning on the dimensions of the data which are orthogonal to $\phi$ is equivalent to conditioning on the values of $W_{1}^{l},\ldots,W_{h-1}^{l}$, $W_{1}^{r},\ldots,W_{h-1}^{r}$, and $C_{0}^{2}$. We can see from (4) above that doing this while allowing only $\phi$ to change has the effect of multiplying $X_{\hat{\tau}-h+1,\ldots,X_{\hat{\tau}}}$ by $\sqrt{\frac{\phi}{\phi_{obs}}}$ and multiplying $X_{\hat{\tau}+1},\ldots,X_{\hat{\tau}+h}$ by $\sqrt{\frac{1-\phi}{1-\phi_{obs}}}$. As we change $\phi$ we are re-scaling the data before and after $\hat{\tau}$ so that the proportion of the sum of squares of the data before and after $\hat{\tau}$ changes whilst the overall sum of squares of the data remains unchanged.
Hence, let
@@
We consider only the subset of possible data sets $\{\boldsymbol{X}^{\prime}(\phi):\phi\in\mathbb{R}\}$. Letting $\mathcal{S}$ denote the set $\mathcal{S}=\{\phi:\hat{\tau}\in\mathcal{M}(\boldsymbol{X}^{\prime}(\phi))\}$ – the set of $\phi$ values such that the changepoint of interest, $\hat{\tau}$, is included as a change in the output of the changepoint algorithm – the post-selection $p$-value we want to calculate is
@@ 3.1 CUSUM for change in variance
$G_{s,e}(t)=\left(\frac{(t-s+1)(e-t)}{e-s+1}\right)^{1/2}\left(\frac{1}{t-s+1}\sum_{j=s}^{t}Y_{j}-\frac{1}{e-t}\sum_{j=t+1}^{e}Y_{j}\right),~{}~{}t=s,\ldots,e-1.$
@@ 3.1.1 Calculating the post-selection $p$-value
To calculate $\mathcal{S}$ we can borrow ideas from methods used for the change in mean model. [20] showed that by conditioning on the set of changes $\mathcal{M}(\boldsymbol{X})$, the order in which the changes are found, and the direction of each change (positive or negative), we can write down a set of inequalities which must be satisfied by the CUSUM statistic. These are derived by considering each iteration of binary segmentation in turn, which is possible as conditioning on the order in which changes are found means we know the segments being tested at each iteration. For each iteration where we detect a change, the CUSUM statistics must be largest at the detected value and be above the threshold, while if no change is detected all the CUSUM statistics are below the threshold. We can write out inequalities for each of these possibilities.
For a given segment $(s,e)$, in order to detect a change point at $\hat{\tau}$ we must have $|G_{s,e}(\hat{\tau})|\geq\lambda$, where $\lambda$ is the changepoint detection threshold, and
@@
$\boldsymbol{X}_{mean}^{\prime}(\phi)=\boldsymbol{X}_{obs}-\frac{\boldsymbol{\nu}}{||\boldsymbol{\nu}||_{2}^{2}}\phi_{obs}+\frac{\boldsymbol{\nu}}{||\boldsymbol{\nu}||_{2}^{2}}\phi.$
Since $G_{s,e}(t)$ is a linear function of $\boldsymbol{X}$, and $\boldsymbol{X}_{mean}^{\prime}(\phi)$ is a linear function of $\phi$, the CUSUM statistic in terms of $\phi$ is also linear in $\phi$. Hence, we can write the set of inequalities in $G_{s,e}(t)$ as linear inequalities in $\phi$, and they will therefore hold for an interval of $\phi$ values, which it is straightforward to calculate. [22] shows that we can increase power by extending this idea to avoid conditioning on the order and directions of the changes, by calculating the intervals associated with each possible combination of changes which yield $\hat{\tau}\in\mathcal{M}(\boldsymbol{X})$. This gives us the set $\mathcal{S}$ which is the union of all such intervals.
@@
$\begin{split}G_{s,e}(t)&=\left(\frac{(t-s+1)(e-t)}{e-s+1}\right)^{1/2}\left(\frac{1}{t-s+1}\sum_{j=s}^{t}Y_{j}-\frac{1}{e-t}\sum_{j=t+1}^{e}Y_{j}\right)\\ &=\left(\frac{(t-s+1)(e-t)}{e-s+1}\right)^{1/2}\left(\frac{1}{t-s+1}\sum_{j=s}^{t}X_{j}^{2}-\frac{1}{e-t}\sum_{j=t+1}^{e}X_{j}^{2}\right),\end{split}$
@@
Using the definition of $\boldsymbol{X}^{\prime}(\phi)$ from (5), we can write ${X_{j}^{\prime}}^{2}(\phi)=a_{j}+b_{j}\phi$ where
$a_{j}=\begin{cases}X_{j}^{2}&j\leq\hat{\tau}-h\text{ or }j>\hat{\tau}+h\\ 0&\hat{\tau}-h<j\leq\hat{\tau}\\ \frac{1}{1-\phi_{obs}}X_{j}^{2}&\hat{\tau}<j\leq\hat{\tau}+h\end{cases}$
$b_{j}=\begin{cases}0&j\leq\hat{\tau}-h\text{ or }j>\hat{\tau}+h\\ \frac{1}{\phi_{obs}}X_{j}^{2}&\hat{\tau}-h<j\leq\hat{\tau}\\ -\frac{1}{1-\phi_{obs}}X_{j}^{2}&\hat{\tau}<j\leq\hat{\tau}+h\end{cases}$
@@ 3.2 Likelihood-ratio based detection of changes
Table 1 demonstrates how this leads to increased accuracy compared with using the CUSUM statistic. We simulated data from a model with $T=400$ and changepoints at $\tau=100,200,300$; the variances on the four segments being $\{1,4,0.25,1\}$, and for each simulated data set we estimated three changes using binary segmentation with both the CUSUM and likelihood ratio statistics. For each true changepoint location, the table records the proportion of runs in which a changepoint was detected within 10 of that location. We see that CUSUM does reasonably well at detecting the first two changes, but does very badly at locating the third, smaller change. Using the likelihood ratio statistic, however, all three changepoints are detected in over $90\%$ of cases.
@@ 3.2.1 Post-selection inference
where $\mathcal{S}=\{\phi:\hat{\tau}\in\mathcal{M}(\boldsymbol{X}^{\prime}(\phi))\}$. Previously, when detecting changes based on the CUSUM statistic, we were able to write down a set of linear inequalities in $\phi$ which defined the set $\mathcal{S}$. However, while we can write down inequalities in $\Lambda_{s,e}(\phi,\tau)$ as we did for $G_{s,e}(\phi,\tau)$, $\Lambda_{s,e}(\phi,\tau)$ is not a linear function of $\phi$, see the Supplementary Material for more details.
@@
One way to reduce the variance of $\hat{p}$, without increasing the computation time, is to use stratified sampling. For $i=1,\ldots,N$, we sample $z^{(i)}\sim U(\frac{i-1}{N},\frac{i}{N})$. We then calculate $\phi^{(i)}$ to satisfy $\Pr(\phi<\phi^{(i)})=z^{(i)}$ (see Section C.2 in the Supplementary Material). However, it is often the case that $\Pr(\phi\in\mathcal{S})$ is of the order of $\frac{1}{N}$ or smaller, in which case using stratified sampling is unlikely to improve the estimate $\hat{p}$, as we will still have very few $\phi^{(i)}\in\mathcal{S}$. To overcome this we resort to using simulation to estimate the intervals that contribute to $\mathcal{S}$.